OpenStudy (anonymous):

Dimension of rectangles are 3x+5 and x+7 Write an expression for the area of the rectangle as a product and as a polynomial in standard form.

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

@321oky ccan u help

OpenStudy (anonymous):

1. 8x6+6x5+2x4 2. 8x3−28x2+4x−6x2+21x−3 =8x3−34x2+25x−3 3. 2x4+x3+6x2+8x3+4x2+24x−6x2−3x−18 =2x4+9x3+4x2+21x−18

OpenStudy (anonymous):

Area of Rectangle = First Side*Other Side or Length*Breadth Therefore, Required Polynomial = 2*(x - 4)*(x + 5) =2(x2+5x−4x−20) =2x2+2x−40

OpenStudy (anonymous):

So what do u right all what u put? on the first comment or second